OpenStudy (anonymous):

If f(x) = x3 – 2x2, which expression is equivalent to f(i)?

OpenStudy (anonymous):

–2 + i –2 – i 2 + i 2 – I My choices

OpenStudy (amoodarya):

\[i=\sqrt{-1}\\i^2=-1\\i^3=i^2i=-i\\i^4=(i^2)^2=(-1)^2=1\\\]

OpenStudy (amoodarya):

\[f(i)=i^3-2i^2=\\i^2i-2(-1)\]

OpenStudy (amoodarya):

can you go on ?

OpenStudy (anonymous):

No, I honestly don't understand this section on complex number I've been working on the same quiz since 8:00

OpenStudy (amoodarya):

you can replace i^3 by -i and i^2=-1 put in function let me see your work

OpenStudy (amoodarya):

where are you ?